# WateRF: Robust Watermarks in Radiance Fields for Protection of Copyrights

### v1

@@ 5.4 Experimental results
Capacity As shown in Tab. 1, there is a clear trade-off between bit accuracy and message bit length. Bit accuracy test is carried out on 4, 8, 16, 32, and 48 bits. Models implemented with our method show a gradual decrease in bit accuracy as the length of the message increases. However, compared to CopyRNeRF [24], the decline in bit accuracy is less dramatic and less steep. Notably, CopyRNeRF [24] shows a modest 60.6% bit accuracy for 48 bits, whereas TensoRF [4] and Vanilla NeRF [29] using our method demonstrates a respectable 85.82% and 70.43% accuracy, respectively. Additionally, to demonstrate the stability of our method across a range of messages, we fine-tune 500 different messages, each corresponding to unique, randomly generated bits. We rigorously evaluate the bit accuracy for all images rendered by TensoRF [4] and Vanilla NeRF [29]. As illustrated in Fig. 3, the plot shows a minor descending trend that corresponds with the increasing number of unique messages, and it shows that our method is robust with using many different messages, which is a key value to identify unique NeRF model.
@@
Invisibility Our method demonstrates a trade-off between bit accuracy and the quality of reconstructed data, as detailed in Tab. 2. This trade-off is analyzed using several metrics: bit accuracy and visual quality metrics. With TensoRF [4], our method achieves the best bit accuracy and PSNR results. Conversely, when applied to Vanilla NeRF [29], it yields the highest SSIM and the lowest LPIPS (noting that higher SSIM and lower LPIPS are desirable). Although the SSIM and LPIPS scores for TensoRF [4] using our method are not the highest, they are comparable to those achieved with Vanilla NeRF [29]. In summary, models trained with our method attain superior bit accuracy, PSNR, SSIM, and the lowest LPIPS, indicating a well-balanced trade-off between bit accuracy and reconstruction quality. Additionally, to verify the invisibility of our Result, We compute the differences between GT and our rendered results. As illustrated in Fig. 4, the differences between the results of applying our method to both TensoRF [4] and Vanilla NeRF [29] are minimal compared to applying it to CopyRNeRF [24]. This indicates that our method can embed watermarks with minimal impact on the image quality.
Robustness We evaluate the robustness of the watermarked image to different attacks by applying Gaussian noise, Rotation, Scaling, Gaussian Blur, Crop, Brightness, JPEG compression and combination of these distortions. As shown in Tab. 3, We experiment on both TensoRF [4] and Vanila NeRF [29] when applied with our method, resulting in robustness against various attacks compared to CopyRNeRF [24].
@@ 5.5 Ablation study
Patch Loss and Frequency Domain This section delves into the effectiveness of integrating patch loss and utilizing the frequency domain. Fig.  5 presents a comparative analysis of our method, applied to Vanilla NeRF [29], both with and without the incorporation of patch loss and use in the frequency domain. As shown in Fig.  5, the absence of DWT significantly decreases bit accuracy. Without the patch-wise loss, there is a considerable decrease in image quality.
Figure 5: Visual quality comparisons We evaluate our full method, our method without patch loss and our method without frequency domain about 16 bits with NeRF [29].
DWT Level During our experiments with Discrete Wavelet Transform (DWT), we evaluate performance across commonly used levels: 1, 2, 3, and 4. As shown in Tab. 4, it is indicated that the results from levels 1, 3, and 4 do not perform as well as the outcomes achieved at level 2, both in terms of bit accuracy and image quality balance. We also conduct applying DFT, DCT, and no frequency instead of DWT. Consequently, our findings show that the DWT at level 2 is the most efficacious. We use the DWT at level 2 based on the results from the ablation experiments.
Figure 6: Bit accuracy comparison between DWT Subbands We evaluate the bit accuracy on level 1,2,3 and all subbands. We show the results on $M_{L}$ = 16 bits. The results are averaged on all examples.
DWT Subband The choice of DWT subband is as critical as the level of DWT itself. Typically, subbands such as HL, LH, and HH, which are composed of high-frequency components less perceptible to the people, are employed for watermark embedding in conventional watermarking techniques. Conversely, in our experimental approach, we opt for the LL subband representing the image’s coarse approximation and containing essential low-frequency details. This decision is based on the watermark decoder’s pre-training to extract watermarks specifically from the image’s spatial domain and its demonstrated robustness to JPEG compression. Referencing Fig. 6, our empirical investigations confirm that the LL subband yields the highest bit accuracy for watermark extraction.
@@
Comparison of Training time. Even though our proposed method necessitates the use of a pre-trained model’s weights as a starting point for fine-tuning, the time requirement for training is significantly less compared to that of CopyRNeRF [24]. As depicted in Fig. 7, it is evident that the model utilizing our techniques substantially outperforms CopyRNeRF [24] regarding training efficiency. Our method, when applied to an explicit NeRF architecture, such as TensoRF [4], achieves a remarkable speed increase, performing up to six times faster than CopyRNeRF [24]. Similarly, implementing our techniques within the Implicit NeRF framework, like Vanilla NeRF [29], yields a significant speed improvement, with our model operating six times faster than CopyRNeRF [24] but it also surpasses the bit accuracy of the compared model.
@@ 6 Conclusion
We propose a neural 3D watermarking method for the NeRF model. Our method trains a 2D watermark decoder and NeRF model separately. Therefore, our pipeline only required training the decoder once and re-use it on different NeRF watermark models. We adopt conventional watermark techniques in image watermarking, which transfer the image from the spatial domain to the frequency domain to encode the watermark into the image efficiently. We find that Discrete Wavelet Transform (DWT) transformation and patch-wise loss enhance the overall image quality while performing high message bit accuracy. Through extensive experiments, we demonstrate that our method outperforms the state-of-the-art method and is six times faster than the previous work.
Limitations. Although our method show outstanding performance in the digital watermark of Neural Radiance Fields, training a decoder that extracts watermark from rendered images is time-consuming. Training the decoder requires approximately 12 hours on a single RTX 3090. Our method does not use a message encoder inside the pipeline; the model only encodes a unique message per model, which requires fine-tuning the NeRF model to encode different messages. We observe that the bit accuracy drops when we apply numerous messages(see Fig. 3).

### v4

@@ 5.4 Experimental results
Capacity As shown in Tab. 1, there is a clear trade-off between bit accuracy and message bit length. We conduct bit accuracy tests on messages of lengths 4, 8, 16, 32, and 48 bits. Models implemented with our method show a gradual decrease in bit accuracy as the length of the message increases. However, when compared to CopyRNeRF [30], the decline in bit accuracy is less dramatic and less steep. Notably, CopyRNeRF [30] shows a modest 60.6% bit accuracy for 48 bits, whereas Vanilla NeRF [35] and TensoRF [3] using our method demonstrates a respectable 85.82% and 70.43% accuracy, respectively. Additionally, to demonstrate the stability of our method across a range of messages, we fine-tune 500 different messages, each corresponding to unique, randomly generated bits. We rigorously evaluate the bit accuracy for all images rendered by Vanilla NeRF [35] and TensoRF [3]. As illustrated in Fig. 3, the plot shows a minor descending trend that corresponds with the increasing number of unique messages, and it shows that our method is robust with using many different messages, which is a key value to identify unique NeRF model.
@@
Invisibility Our method demonstrates a trade-off between bit accuracy and the quality of reconstructed data, as detailed in Tab. 2. This trade-off is analyzed using several metrics, including bit accuracy and visual quality metrics. With TensoRF [3], our method achieves the best bit accuracy and PSNR results. Conversely, when applied to Vanilla NeRF [35], it yields the highest SSIM and the lowest LPIPS (noting that higher SSIM and lower LPIPS are desirable). Although the SSIM and LPIPS scores for TensoRF [3] using our method are not the highest, they are comparable to those achieved with Vanilla NeRF [35]. In summary, models trained with our method attain superior bit accuracy, PSNR, SSIM, and the lowest LPIPS, indicating a well-balanced trade-off between bit accuracy and reconstruction quality.
Additionally, to verify the invisibility of our Result, We compute the differences between ground truth and our rendered results. As illustrated in Fig. 4, the results of applying our method to both Vanilla NeRF [35] and TensoRF [3] achieve a better balance between reconstruction quality and bit accuracy compared to CopyRNeRF [30]. This indicates that our method can embed watermarks with minimal impact on the reconstruction quality.
Robustness We evaluate the robustness of the watermarked image to different attacks by applying Gaussian noise, Rotation, Scaling, Gaussian Blur, Crop, Brightness, JPEG compression and combination of these distortions. As shown in Tab. 3, We experiment on both Vanila NeRF [35] and TensoRF [3] when applied with our method, resulting in robustness against various attacks compared to CopyRNeRF [30].
@@ 5.5 Ablation study
Patch Loss and Frequency Domain This section delves into the effectiveness of integrating patch loss and utilizing the frequency domain. Fig.  5 presents a comparative analysis of our method, applied to Vanilla NeRF [35], both with and without the incorporation of patch loss and use in the frequency domain. As shown in Fig.  5, the absence of DWT significantly decreases bit accuracy. Without the patch-wise loss, there is a decrease in reconstruction quality.
Figure 5: Reconstruction quality comparisons We evaluate our full method, our method without patch loss and our method without frequency domain about 16 bits with NeRF [35].
DWT Level During our experiments with DWT, we evaluate performance across commonly used levels: 1, 2, 3, and 4. As shown in Tab. 4, it is indicated that the results from levels 1, 3, and 4 do not perform as well as the outcomes achieved at level 2, both in terms of bit accuracy and reconstruction quality balance. Additionally, we conduct applying DFT, DCT, and no frequency instead of DWT. As a result, our findings show that the DWT at level 2 is the most effective. Thus, we use the DWT at level 2 based on the results from the ablation experiments.
@@
DWT Subband The choice of DWT subband is as critical as the level of DWT itself. Typically, subbands such as LH, HL, and HH, which are composed of high-frequency components less perceptible to the people, are employed for watermark embedding in conventional watermarking techniques. Conversely, in our experimental approach, we opt for the LL subband representing the image’s coarse approximation and containing essential low-frequency details. This decision is based on the watermark decoder’s pre-training process to extract watermarks specifically from the image’s spatial domain, which is visually similar to the LL subband and its demonstrated robustness to JPEG compression. In Fig. 6, our empirical investigations confirm that the LL subband yields the highest bit accuracy.
Figure 6: Bit accuracy comparison between DWT Subbands We evaluate the bit accuracy on level 1,2,3 and all subbands. We show the results on $M_{L}$ = 16 bits.
Comparison of Training time. Even though our proposed method necessitates the use of a pre-trained model’s weights as a starting point for fine-tuning NeRF, the time requirement for training is significantly less compared to that of CopyRNeRF [30]. As depicted in Fig. 7, it is evident that the model utilizing our techniques substantially outperforms CopyRNeRF [30] regarding training efficiency. Our method, when applied to an explicit NeRF architecture, such as TensoRF [3], achieves a remarkable speed increase, performing up to six times faster than CopyRNeRF [30]. Similarly, implementing our techniques within the Implicit NeRF framework, like Vanilla NeRF [35], yields a significant speed improvement, with our model operating six times faster than CopyRNeRF [30] but it also surpasses the bit accuracy of the compared model.
@@ 6 Conclusion
We propose a neural 3D watermarking method for the NeRF model. Our method trains a 2D watermark decoder and NeRF model separately. Therefore, our pipeline only requires training the decoder once and re-use it on different NeRF watermark models. We adopt conventional watermark techniques in image watermarking, which transfer the image from the spatial domain to the frequency domain to encode the watermark into the image efficiently. We find that DWT and patch-wise loss enhance the overall reconstruction quality while performing high message bit accuracy. Through extensive experiments, we demonstrate that our method outperforms the state-of-the-art method and is six times faster than the previous work.
Limitations. Our method shows outstanding performance in the digital watermark of NeRF, but training a decoder that extracts watermark from rendered images is time-consuming, taking approximately 12 hours on a single RTX 3090. Our method only encodes a unique message per model, requiring fine-tuning for message insertion every time. We observe that the bit accuracy drops when we apply numerous messages(see Fig. 3). Although previous studies, including us, report robustness against various distortions of images, there is no consideration about cases where the model is attacked. Future work may explore watermarks resilient to cases of model attacks.
Acknowledgments. This research was supported by the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture((Project: 4D Content Generation and Copyright Protection with Artificial Intelligence, R2022020068, 80%),(Project: International Collaborative Research and Global Talent Development for the Development of Copyright Management and Protection Technologies for Generative AI, RS-2024-00345025, 10%)), the National Research Foundation of Korea grant (NRF-2022R1F1A1074334, 5%), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT)(No. 2019-0-00079, Artificial Intelligence Graduate School Program(Korea University), 5%).